# Tokenization Is More Than Compression

### v1

@@ Abstract
Craig W. Schmidt  Varshini Reddy  Haoran Zhang†,‡  Alec Alameddine Omri Uzan§  Yuval Pinter§  Chris Tanner†,¶ Kensho Technologies Harvard University §Department of Computer Science MIT Cambridge, MA Cambridge, MA Ben-Gurion University of the Negev Cambridge, MA Beer Sheva, Israel {craig.schmidt,varshini.reddy,alec.alameddine,chris.tanner}@kensho.com haoran_zhang@g.harvard.edu{omriuz@post,uvp@cs}.bgu.ac.il
@@ 1 Introduction
Tokenization is an essential step in NLP that translates human-readable sentences into a sequence of distinct tokens that can be subsequently used by statistical models Grefenstette (1999). Recently, a growing number of studies have researched the effects of tokenization, both in an intrinsic manner and as it affects downstream model performance Singh et al. (2019); Bostrom and Durrett (2020); Hofmann et al. (2021, 2022); Limisiewicz et al. (2023). To rigorously inspect the impact of tokenization, we consider tokenization as three distinct, sequential stages:
@@
    Segmentation: given a vocabulary $\mathcal{V}$ and a document $d$, segmentation determines how to split $d$ into a series of $K_{d}$ tokens $t_{1},\dots,t_{k},\dots,t_{K_{d}}$, with all $t_{k}\in\mathcal{V}$, such that the concatenation of the tokens strictly equals $d$. Given a corpus of documents $d\in\mathcal{C}$, we will define the corpus token count (CTC) as the total number of tokens used in each segmentation, $CTC=\sum_{d\in\mathcal{C}}K_{d}$.
@@
The widely used Byte-Pair Encoding (BPE) tokenizer Sennrich et al. (2016) originated in the field of data compression Gage (1994). Gallé (2019) argues that it is effective because it compresses text to a short sequence of tokens. To investigate the hypothesis that having fewer tokens leads to better downstream performance, we design a novel tokenizer, PathPiece, that, for a given document $d$ and vocabulary $\mathcal{V}$, finds a segmentation with the minimum possible $K_{d}$. The PathPiece vocabulary construction routine is a top-down procedure that heuristically minimizes CTC on a training corpus. PathPiece is ideal for studying the effect of CTC on downstream performance, as we can vary decisions at each tokenization stage.
We extend these experiments to the most commonly used tokenizers, focusing on how downstream task performance is impacted by the major stages of tokenization and vocabulary sizes. Toward this aim, we conducted experiments by training 64 language models (LMs): 54 LMs with 350M parameters; 6 LMs with 1.3B parameters; and 4 LMs with 2.4B parameters. We provide open-source, public access to PathPiece, our token vocabularies, and all our 64 trained LMs. [^1] https://github.com/kensho-technologies
2 Related Work
@@ 2.1 Pre-tokenization Methods
Pre-tokenization is a process of breaking text into chunks, which are then tokenized independently. A token is not allowed to cross these pre-tokenization boundaries. BPE, WordPiece, and Unigram all require new chunks to begin whenever a space is encountered. If a space appears in a chunk, it must be the first character; hence, we will call this “FirstSpace”. Thus “␣New” is allowed but “New␣York” is not. Gow-Smith et al. (2022) suggest treating spaces as individual tokens, which we will call “Space” pre-tokenization, while Jacobs and Pinter (2022) suggest leaving spaces at the end of tokens. Llama Touvron et al. (2023) popularized the idea of having each digit always be an individual token, which we call “Digit” pre-tokenization.
@@ 2.2 Vocabulary Construction
We focus on byte-level, lossless subword tokenization. Subword tokenization algorithms split text into word and subword units based on their frequency and co-occurrence patterns from their “training” data, effectively capturing morphological and semantic nuances in the tokenization process Mikolov et al. (2011). Mielke et al. (2021) gives a survey of subword tokenization.
We analyze BPE, WordPiece, and Unigram as baseline subword tokenizers, using the implementations from HuggingFace [^2] https://github.com/huggingface/tokenizers with ByteLevel pre-tokenization enabled. We additionally study SaGe, a context-sensitive subword tokenizer, using the recently released 2.0 version [^3] https://github.com/MeLeLBGU/SaGe.
@@
Sennrich et al. (2016) introduced Byte-Pair Encoding (BPE), a bottom-up method where the vocabulary construction starts with single bytes as tokens. It then merges the most commonly occurring pair of adjacent tokens in a training corpus into a single new token in the vocabulary. This process repeats until the desired vocabulary size is reached. Issues with BPE and analyses of its properties are discussed in Bostrom and Durrett (2020); Klein and Tsarfaty (2020); Gutierrez-Vasques et al. (2021); Yehezkel and Pinter (2023); Saleva and Lignos (2023); Liang et al. (2023).
@@ 2.3 Segmentation Methods
Certain segmentation methods are tightly coupled to the vocabulary construction step, such as merge rules for BPE or the maximum likelihood approach for Unigram. Others, such as the WordPiece approach of greedily taking the longest prefix token in the vocabulary at each point, can be applied to any vocabulary. Alternative segmentation schemes include Dynamic Programming BPE He et al. (2020), BPE-Dropout Provilkov et al. (2020), and FLOTA Hofmann et al. (2022).
@@ 3 PathPiece
Several efforts over the last few years (Gallé, 2019; Zouhar et al., 2023, inter alia) have suggested that the empirical advantage of BPE as a tokenizer in many NLP applications, despite its unawareness of language structure, can be traced to its superior compression abilities, providing models with overall shorter sequences during learning and inference. Inspired by this claim we introduce PathPiece, a lossless subword tokenizer that, given a vocabulary $\mathcal{V}$ and document $d$, produces a segmentation minimizing the total number of tokens needed to split $d$. We additionally provide a vocabulary construction procedure that, using this segmentation, attempts to find a $\mathcal{V}$ minimizing the corpus token count (CTC). [^4] An extended description is given in Appendix A. PathPiece provides an ideal testing laboratory for the compression hypothesis by virtue of its maximally efficient segmentation.
@@ 3.1 Segmentation
PathPiece requires that all single-byte tokens are included in vocabulary $\mathcal{V}$ to run correctly. PathPiece works by finding a shortest path through a directed acyclic graph (DAG), where each byte $i$ of training data forms a node in the graph, and two nodes $j$ and $i$, contain a directed edge if the byte segment $[j,i]$ is a token in $\mathcal{V}$.
We describe PathPiece segmentation in Algorithm 1, where $L$ is a limit on the maximum width of a token in bytes, which we set to 16. It has a complexity of $O(nL)$, which follows directly from the two nested for-loops. For each byte $i$ in $d$, it computes the shortest path length $pl[i]$ in tokens up to and including byte $i$, and the width $wid[i]$ of a token with that shortest path length. In choosing $wid[i]$, ties between multiple tokens with the same shortest path length $pl[i]$ can be broken randomly, or the one with the longest $wid[i]$ can be chosen, as shown here. [^5] Random tie-breaking, which can be viewed as a form of subword regularization, is presented in Appendix A. Some motivation for selecting the longest token is due to the success of FLOTA Hofmann et al. (2022). Then, a backward pass constructs the shortest possible segmentation from the $wid[i]$ values computed in the forward pass.
@@ 5 Results
Table 1: Summary of 350M parameter model downstream accuracy (%). The Overall Avg column averages across the three vocabulary sizes. The Rank column refers to the Overall Avg column, best to worst.
@@ 5.1 Vocabulary Size
Figure 1 gives the overall average, along with the individual averages, for each of the three vocabulary sizes for each variant, labeled according to the rank from Table 1. We observe that there is a high correlation between downstream performance at different vocabulary sizes. The pairwise $R^{2}$ values for the accuracy of the 32,768 and 40,960 runs was 0.750; between 40,960 and 49,152 it was 0.801; and between 32,768 and 49,152 it was 0.834. This indicates that vocabulary size is not a crucial decision over this range of sizes. Given this high degree of correlation, we focus our analysis on the overall average accuracy. This averaging removes some of the variance amongst individual language model runs. Thus, unless specified otherwise, our analyses present performance averaged over vocabulary sizes.
@@ 5.2 Overall performance
Figure 2: Pairwise $p$-values for 350M model results. Black boxes lines represent $p$ > 0.05. The top 6 tokenizers are all competitive, and there is no statistically significantly best approach.
@@
All the results reported thus far are for language models with identical architectures and 350M parameters. To examine the dependency on model size, we build larger models of 1.3B parameters for six of our experiments, and 2.4B parameters for four of them. In the interest of computational time, these larger models were only run at a single vocabulary size of 40,960. The average results over the 10 task accuracies for these models are given in Figure 6 in Appendix D. See Figure 7 in Appendix D for an example checkpoint graph at each model size. The main result from these models is that the relative performance of the tokenizers does vary by model size, and that there is a group of highly-performant tokenizers that yield comparable results.
This aligns with our finding that the top six tokenizers are not statistically better than one another at the 350M model size.
@@ 5.3 Corpus Token Count vs Accuracy
Figure 3 shows the corpus token count (CTC) versus the accuracy of each vocabulary size. We do not find a straightforward relationship between the two.
@@
The two models with the highest CTC are: PathPiece with Space pre-tokenization (12), which is to be expected given each space is its own token; and, SaGe with an initial Unigram vocabulary (10) also had a high CTC. In terms of accuracy, PathPiece with no pre-tokenization (18) and Unigram with PathPiece segmentation (17) both did quite poorly. Notably, the range of CTC is quite narrow within each vocabulary construction method, even while changes in pre-tokenization and segmentation lead to significant accuracy differences. While there are confounding factors present in this chart (e.g., pre-tokenization, vocabulary initialization, and that more tokens allow for additional computations by the downstream model) it is difficult to discern any trend that lower CTC leads to improved performance. If anything, there seems to be an inverted U-shaped curve with respect to the CTC and downstream performance.
@@ 6 Analysis
We now analyze the results across the various experiments in a more controlled manner. Our experiments allow us to examine changes in each stage of tokenization, holding the rest constant, revealing design decisions making a significant difference.
@@ 6.1 Pre-tokenization
For PathPieceR with an $n$-gram initial vocabulary, we can isolate pre-tokenization. PathPiece is efficient enough to process entire documents with no pre-tokenization, giving it full freedom to minimize the corpus token count (CTC). Adding pre-tokenization constrains PathPiece’s ability to minimize tokens, giving a natural way to vary the number of tokens. [^11] These two runs also used Digit pre-tokenization where each digit is its own token.
Figure 4 shows that PathPiece, as intended, minimizes the number of tokens used over a corpus when trained with no pre-tokenization (18). The other variants restrict spaces to either be the first character of a token (14), or their own token (12). Gallé (2019) suggests we should expect downstream performance to decrease with more tokens. Interestingly, however, we observe an increase in overall accuracy for our downstream tasks, as a function of the number of tokens. These results suggest it is not inherently beneficial to use fewer tokens, aligning with our findings in §5.3.
@@ 18 None The ␣valu ation␣is ␣estimated ␣to␣b e␣\$ 2 1 3 M
To understand why, consider the example PathPiece tokenization in Table 2 for the three pre-tokenization methods of Figure 4. The None mode uses the word-boundary-spanning tokens “ation␣is”, “␣to␣b”, and “e␣\$”. The lack of morphological alignment demonstrated in this example is likely more important to downstream model performance than a simple token count.
Considering the $p$-values of Figure 2, SpaceDigit (12) significantly improves accuracy over FirstSpaceDigit (14) and No pre-tokenization (18). FirstSpaceDigit (14) also performed significantly better than No pre-tokenization (18).
One of our examples that isolated Digit pre-tokenization showed it was worse with significance (8 versus 11), while the other (15 versus 16) was not statistically significant. Other than mathqa, none of our downstream tasks were particularly mathematical in nature, making it difficult to conclude a definitive judgment about the utility of Digit pre-tokenization. [^12] See Figure 8 and Figure 9 in Appendix E for more details.

### v2

@@ Abstract
Craig W. Schmidt  Varshini Reddy  Haoran Zhang†,‡  Alec Alameddine Omri Uzan§  Yuval Pinter§  Chris Tanner†,¶ Kensho Technologies Harvard Univ §Ben-Gurion University MIT Cambridge, MA Cambridge, MA Beer Sheva, Israel Cambridge, MA {craig.schmidt,varshini.reddy,alec.alameddine,chris.tanner}@kensho.com haoran_zhang@g.harvard.edu{omriuz@post,uvp@cs}.bgu.ac.il
@@ 1 Introduction
Tokenization is an essential step in NLP that translates human-readable text into a sequence of distinct tokens that can be subsequently used by statistical models Grefenstette (1999). Recently, a growing number of studies have researched the effects of tokenization, both in an intrinsic manner and as it affects downstream model performance Singh et al. (2019); Bostrom and Durrett (2020); Hofmann et al. (2021, 2022); Limisiewicz et al. (2023); Zouhar et al. (2023b). To rigorously inspect the impact of tokenization, we consider tokenization as three distinct, sequential stages:
@@
    Segmentation: given a vocabulary $\mathcal{V}$ and a document $d$, segmentation determines how to split $d$ into a series of $K_{d}$ tokens $t_{1},\dots,t_{k},\dots,t_{K_{d}}$, with all $t_{k}\in\mathcal{V}$, such that the concatenation of the tokens strictly equals $d$. Given a corpus of documents $\mathcal{C}$, we will define the corpus token count (CTC) as the total number of tokens used in each segmentation, $\operatorname{CTC}(\mathcal{C})=\sum_{d\in\mathcal{C}}K_{d}$.
@@
The widely used Byte-Pair Encoding (BPE) tokenizer Sennrich et al. (2016) originated in the field of data compression Gage (1994). Gallé (2019) argues that it is effective because it compresses text to a short sequence of tokens. Goldman et al. (2024) varied the number of documents in the tokenizer training data for BPE, and found a correlation between CTC and downstream performance. To investigate the hypothesis that having fewer tokens necessarily leads to better downstream performance, we design a novel tokenizer, PathPiece, that, for a given document $d$ and vocabulary $\mathcal{V}$, finds a segmentation with the minimum possible $K_{d}$. The PathPiece vocabulary construction routine is a top-down procedure that heuristically minimizes CTC on a training corpus. PathPiece is ideal for studying the effect of CTC on downstream performance, as we can vary decisions at each tokenization stage.
We extend these experiments to the most commonly used tokenizers, focusing on how downstream task performance is impacted by the major stages of tokenization and vocabulary sizes. Toward this aim, we conducted experiments by training 64 language models (LMs): 54 LMs with 350M parameters; 6 LMs with 1.3B parameters; and 4 LMs with 2.4B parameters. We provide open-source, public access to PathPiece, [^1] https://github.com/kensho-technologies/pathpiece and our trained vocabularies and LMs. [^2] https://github.com/kensho-technologies/timtc_vocabs_models
2 Preliminaries
Ali et al. (2024) and Goldman et al. (2024) examined the effect of tokenization on downstream performance of LLM tasks, reaching opposite conclusions on the importance of CTC. Zouhar et al. (2023a) also find that low token count alone does not necessarily improve performance. Mielke et al. (2021) give a survey of subword tokenization.
@@ 2.1 Pre-tokenization Methods
Pre-tokenization is a process of breaking text into chunks, which are then tokenized independently. A token is not allowed to cross these pre-tokenization boundaries. BPE, WordPiece, and Unigram all require new chunks to begin whenever a space is encountered. If a space appears in a chunk, it must be the first character; hence, we will call this “FirstSpace”. Thus “␣New” is allowed but “New␣York” is not. Gow-Smith et al. (2022) examine treating spaces as individual tokens, which we will call “Space” pre-tokenization, while Jacobs and Pinter (2022) suggest marking spaces at the end of tokens, and Gow-Smith et al. (2024) propose dispensing them altogether in some settings. Llama Touvron et al. (2023) popularized the idea of having each digit always be an individual token, which we call “Digit” pre-tokenization.
@@ 2.2 Vocabulary Construction
We focus on byte-level, lossless subword tokenization. Subword tokenization algorithms split text into word and subword units based on their frequency and co-occurrence patterns from their “training” data, effectively capturing morphological and semantic nuances in the tokenization process Mikolov et al. (2011).
We analyze BPE, WordPiece, and Unigram as baseline subword tokenizers, using the implementations from HuggingFace [^3] https://github.com/huggingface/tokenizers with ByteLevel pre-tokenization enabled. We additionally study SaGe, a context-sensitive subword tokenizer, using version 2.0.444https://github.com/MeLeLBGU/SaGe
@@
[…]
@@ 2.3 Segmentation Methods
Certain segmentation methods are tightly coupled to the vocabulary construction step, such as merge rules for BPE or the maximum likelihood approach for Unigram. Others, such as the WordPiece approach of greedily taking the longest prefix token in the vocabulary at each point, can be applied to any vocabulary; indeed, there is no guarantee that a vocabulary will perform best downstream with the segmentation method used to train it Uzan et al. (2024). Additional segmentation schemes include Dynamic Programming BPE He et al. (2020), BPE-Dropout Provilkov et al. (2020), and FLOTA Hofmann et al. (2022).
@@ 3 PathPiece
Several efforts over the last few years (Gallé, 2019; Zouhar et al., 2023a, inter alia) have suggested that the empirical advantage of BPE as a tokenizer in many NLP applications, despite its unawareness of language structure, can be traced to its superior compression abilities, providing models with overall shorter sequences during learning and inference. Inspired by this claim we introduce PathPiece, a lossless subword tokenizer that, given a vocabulary $\mathcal{V}$ and document $d$, produces a segmentation minimizing the total number of tokens needed to split $d$. We additionally provide a vocabulary construction procedure that, using this segmentation, attempts to find a $\mathcal{V}$ minimizing the corpus token count (CTC). [^5] An extended description is given in Appendix A. PathPiece provides an ideal testing laboratory for the compression hypothesis by virtue of its maximally efficient segmentation.
@@ 3.1 Segmentation
PathPiece requires that all single-byte tokens are included in vocabulary $\mathcal{V}$ to run correctly. PathPiece works by finding a shortest path through a directed acyclic graph (DAG), where each byte $i$ of training data forms a node in the graph, and two nodes $j$ and $i$ contain a directed edge if the byte segment $[j,i]$ is a token in $\mathcal{V}$. We describe PathPiece segmentation in Algorithm 1, where $L$ is a limit on the maximum width of a token in bytes, which we set to 16. It has a complexity of $O(nL)$, which follows directly from the two nested for-loops. For each byte $i$ in $d$, it computes the shortest path length $pl[i]$ in tokens up to and including byte $i$, and the width $wid[i]$ of a token with that shortest path length. In choosing $wid[i]$, ties between multiple tokens with the same shortest path length $pl[i]$ can be broken randomly, or the one with the longest $wid[i]$ can be chosen, as shown here. [^6] Random tie-breaking, which can be viewed as a form of subword regularization, is presented in Appendix A. Some motivation for selecting the longest token is due to the success of FLOTA Hofmann et al. (2022). Then, a backward pass constructs the shortest possible segmentation from the $wid[i]$ values computed in the forward pass.
@@ 5 Results
Rank Vocab Constr Init Voc Pre-tok Segment Overall 32,768 40,960 49,152
@@
[…]
@@ 5.1 Vocabulary Size
Figure 1 gives the overall average, along with the individual averages, for each of the three vocabulary sizes for each variant, labeled according to the rank from Table 1. We observe that there is a high correlation between downstream performance at different vocabulary sizes. The pairwise $R^{2}$ values for the accuracy of the 32,768 and 40,960 runs was 0.750; between 40,960 and 49,152 it was 0.801; and between 32,768 and 49,152 it was 0.834. This corroborates the effect shown graphically in Figure 1 that vocabulary size is not a crucial decision over this range of sizes. Given this high degree of correlation, we focus our analysis on the overall average accuracy. This averaging removes some of the variance amongst individual language model runs. Thus, unless specified otherwise, our analyses present performance averaged over vocabulary sizes.
@@ 5.2 Overall performance
Figure 2: Pairwise $p$-values for 350M model results. Boxes outlined in black represent $p$ > 0.05. The top 6 tokenizers are all competitive, and there is no statistically significantly best approach.
@@
All the results reported thus far are for language models with identical architectures and 350M parameters. To examine the dependency on model size, we trained larger models of 1.3B parameters for six of our experiments, and 2.4B parameters for four of them. In the interest of computational time, these larger models were only trained with a single vocabulary size of 40,960. In Figure 6 in subsection 6.4, we report models’ average performance across 10 tasks. See Figure 7 in Appendix D for an example checkpoint graph at each model size. The main result from these models is that the relative performance of the tokenizers does vary by model size, and that there is a group of high performing tokenizers that yield comparable results. This aligns with our finding that the top six tokenizers are not statistically better than one another at the 350M model size.
@@ 5.3 Corpus Token Count vs Accuracy
Figure 3 shows the corpus token count (CTC) versus the accuracy of each vocabulary size, given in Table 11. We do not find a straightforward relationship between the two. Ali et al. (2024) recently examined the relationship between CTC and downstream performance for three different tokenizers, and also found it was not correlated on English language tasks.
@@
The two models with the highest CTC are PathPiece with Space pre-tokenization (12), which is to be expected given each space is its own token, and SaGe with an initial Unigram vocabulary (10). The Huggingface Unigram models in Figure 3 had significantly higher CTC than the corresponding BPE models, unlike Bostrom and Durrett (2020) and Gow-Smith et al. (2022), which report a difference of only a few percent with SentencePiece Unigram. Ali et al. (2024) point out that due to differences in pre-processing, the Huggingface Unigram tokenizer behaves quite differently than the SentencePiece Unigram tokenizer, which may explain this discrepancy.
In terms of accuracy, PathPiece with no pre-tokenization (18) and Unigram with PathPiece segmentation (17) both did quite poorly. Notably, the range of CTC is quite narrow within each vocabulary construction method, even while changes in pre-tokenization and segmentation lead to significant accuracy differences. While there are confounding factors present in this chart (e.g., pre-tokenization, vocabulary initialization, and that more tokens allow for additional computations by the downstream model) it is difficult to discern any trend that lower CTC leads to improved performance. If anything, there seems to be an inverted U-shaped curve with respect to the CTC and downstream performance. The Pearson correlation coefficient between CTC and average accuracy was found to be 0.241. Given that a lower CTC value signifies greater compression, this result suggests a weak negative relationship between the amount of compression and average accuracy.
Zouhar et al. (2023a) introduced an information-theoretic measure based on Rényi efficiency that correlates with downstream performance for their application. [^12] Except, so far, for a family of adversarially-created tokenizers Cognetta et al. (2024). It has an order parameter $\alpha$, with a recommended value of 2.5. We present the Rényi efficiencies and CTC for all models in Table 11 in Appendix G, and summarize their Pearson correlation with average accuracy in Table 2. For the data of Figure 3, all the correlations for various $\alpha$ also have a weak negative association. They are slightly less negative than the association for CTC, although it is not nearly as large as the benefit they saw over sequence length in their application. We note the strong relationship between compression and Rényi efficiency, as the Pearson correlation of CTC and Rényi efficiency with $\alpha$=2.5 is --0.891.
Comparison Pearson Correlation
CTC and Ave Acc 0.241
Rényi Eff and Ave Acc ($\alpha$=1.5) --0.221
Rényi Eff and Ave Acc ($\alpha$=2.0) --0.169
Rényi Eff and Ave Acc ($\alpha$=2.5) --0.151
Rényi Eff and Ave Acc ($\alpha$=3.0) --0.144
Rényi Eff and Ave Acc ($\alpha$=3.5) --0.141
CTC and Rényi Eff ($\alpha$=2.5) --0.891
Table 2: Pearson Correlation of CTC and Average Accuracy, or Rényi efficiency for various orders $\alpha$ with Average Accuracy, or CTC and Rényi efficiency at $\alpha=2.5$.
By varying aspects of BPE, Gallé (2019) and Goldman et al. (2024) suggests we should expect downstream performance to decrease with CTC, while in contrast Ali et al. (2024) did not find a strong relation when varying the tokenizer. Our extensive results varying a number of stages of tokenization suggest it is not inherently beneficial to use fewer tokens. Rather, the particular way that the CTC is varied can lead to different conclusions.
@@ 6 Analysis
We now analyze the results across the various experiments in a more controlled manner. Our experiments allow us to examine changes in each stage of tokenization, holding the rest constant, revealing design decisions making a significant difference. [^13] Appendix E contains additional analysis
@@ 6.1 Pre-tokenization
For PathPieceR with an $n$-gram initial vocabulary, we can isolate pre-tokenization. PathPiece is efficient enough to process entire documents with no pre-tokenization, giving it full freedom to minimize the corpus token count (CTC).
@@ 18 None The ␣valu ation␣is ␣estimated ␣to␣b e␣\$ 2 1 3 M
Adding pre-tokenization constrains PathPiece’s ability to minimize tokens, giving a natural way to vary the number of tokens. Figure 4 shows that PathPiece minimizes the number of tokens used over a corpus when trained with no pre-tokenization (18). The other variants restrict spaces to either be the first character of a token (14), or their own token (12). [^14] These two runs also used Digit pre-tokenization where each digit is its own token. Consider the example PathPiece tokenization in Table 3 for the three pre-tokenization methods. The None mode uses the word-boundary-spanning tokens “ation␣is”, “␣to␣b”, and “e␣\$”. The lack of morphological alignment demonstrated in this example is likely more important to downstream model performance than a simple token count.
In Figure 4 we observe a statistically significant increase in overall accuracy for our downstream tasks, as a function of CTC. Gow-Smith et al. (2022) found that Space pre-tokenization lead to worse performance, while removing the spaces entirely helps [^15] Although omitting the spaces entirely does not lead to a reversible tokenization as we have been considering.. Thus, this particular result may be specific to PathPieceR.